Becky is getting ready to do her taxes. She is single and lives in Denver. Becky earned $110,000 in taxable income in 2015. She reviews the following table, which shows the IRS tax rates for a single taxpayer in 2015.

On Annual Taxable Income...The Tax Rate Is... (Percent)
Up to $9,22510.0
From $9,225 to $37,45015.0
From $37,450 to $90,75025.0
From $90,750 to $189,30028.0
From $189,300 to $411,50033.0
From $411,500 to $413,20035.0
Over $413,20039.6


Based on the IRS table, Becky calculates that her marginal tax rate is _______  when her annual taxable income is $110,000. 

Becky calculates that she owes _______  in income taxes for 2015 .

Becky then calculates that her average tax rate is _______  , based on the annual income level and the amount of taxes she owes for 2015 

After figuring out what she owes in taxes in 2015, Becky decides to ask an accountant for tax advice. The accountant claims that he has found a legal way to shelter $2,000 of taxable income from the federal government. 

The maximum amount that Becky is willing to pay to learn this strategy and reduce her taxable income by $2,000 is _______ .(Hint: Shelterinsome income means finding a legal way to avoid being charged income tax on that income. For example, someone who has $50,000 in taxable income and shelters $10,000 pays income tax on only $40,000.) g
Answer #1

Marginal tax rate= tax rate on last dollar of income = 28%


Total tax = 0.1*9225+0.15*(37450-9225)+0.25*(90750-37450)+0.28*(110000-90750) = $23871.25


Average tax rate = total tax/total income = 23871.25/110000 = 21.70%


Maximum Becky willing to pay = Total tax before sheltering-total tax after sheltering = 0.28*2000 = $560 Because marginal tax rate is 28% and taxable income will decrease by 2000, so tax change will be 0.28*2000


answered by: gavin